OpenStudy (anonymous):

What is the solution set of |2x – 3| – 6 = –1?

OpenStudy (anonymous):

Start with moving -6 to the right side, can you?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

|2x-3|=5

OpenStudy (anonymous):

Now in order to remove the absolute value sign, you'll solve 2 cases: 2x -3 = 5 or 2x - 3 = -5 Can you try?

OpenStudy (anonymous):

2x=8 or 2x=-2

OpenStudy (anonymous):

divide now?

OpenStudy (anonymous):

yes by 2

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

What's your answer?

OpenStudy (anonymous):

4 or -1

OpenStudy (anonymous):

Correct! Now you know the method to solve for absolute value equation?

OpenStudy (anonymous):

yup:) thanks for your help